OpenStudy (anonymous):

A mortgage has the terms shown below: Balloon mortgage $400,000 20/6 balloon 8% annual rate Initial Monthly Payment: $3,345.76 What is the amount of the balloon payment rounded to the nearest dollar?

OpenStudy (anonymous):

The choices are $377,803 $339,756 $337,506 $$159,105

OpenStudy (anonymous):

No matter what I do I keep getting $340,851.99, which isn't one of the choices. Someone please tell me what I'm doing wrong, and walk me through it. Thank you. I appreciate it greatly.

OpenStudy (anonymous):

calculate the monthly payment for a 20 fixed loan at 8% then use that to balance out the loan after 6 years

OpenStudy (anonymous):

P = 400000k^n(1-k)/(1-k^n) n = 20*12 = 240 k = 1+.08/12 400000k^(240)(1-k)/(1-k^(240)); k = 1+.08/12 what is your monthly payment you got in your calcs?

OpenStudy (anonymous):

$3,345.76

OpenStudy (anonymous):

lol, i see they give it too you already ... so much for reading ALL of the question: http://www.wolframalpha.com/input/?i=400000k%5E%28240%29%281-k%29%2F%281-k%5E%28240%29%29%3B+k+%3D+1%2B.08%2F12 3345.76 is what i get too

OpenStudy (anonymous):

so to balance it out ... A = Bk^n + P(1-k^n)/(1-k) B = -400,000, P is given, k is from before, and n=6*12 this time = 72

OpenStudy (anonymous):

Even after getting that, when plugging it into a calculator, I still get the monthly payment I mentioned above, which isn't one of the choices. I don't understand what I'm doing wrong.

OpenStudy (anonymous):

how would you find the balance of a fixed rate mortagage after 6 years?

OpenStudy (anonymous):

im getting one of the answers but i use my own formula for it

OpenStudy (anonymous):

What formula do you use?

OpenStudy (anonymous):

\[A=Bk^n+P\frac{1-k^n}{1-k}\] \[B=initial,~P=Payment,~k=compunded~rate, ~n=periods\]

OpenStudy (anonymous):

as such\[\Large A=(-400000)k^{6*12}+P\frac{1-k^{6*12}}{1-k}\]

OpenStudy (anonymous):

i got no idea how to plug that into a financial calculator tho

OpenStudy (anonymous):

I plugged it in and got $337,506.

OpenStudy (anonymous):

me too

OpenStudy (anonymous):

Thank you for your help. That was really tough. Now it's onward for me to even more difficult questions.

OpenStudy (anonymous):

good luck :)